OpenStudy (zenmo):

(A) Find a set of parametric equations and (B) symmetric equations that passes through the given two points of a line. (-1/2, 2, 1/2), (1, -1/2, 0)

OpenStudy (michele_laino):

how are defined symmetric equations?

OpenStudy (dan815):

probably just meants change them into z=f(x,y)

OpenStudy (michele_laino):

so we have to eliminate the parameter t, right?

OpenStudy (michele_laino):

please keep in mind that a line, in the euclidean space, is given as an intersection between two planes

OpenStudy (michele_laino):

so, all what we can do, is to solve the third equation for t, and then substitute that value of t into the first and second equation

OpenStudy (michele_laino):

namely: \[\Large t = \frac{1}{2} - z\]

OpenStudy (michele_laino):

I got: \[\begin{gathered} x = 1 - 3z \hfill \\ 2y = - 1 + 10z \hfill \\ \end{gathered} \]

OpenStudy (dan815):